• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 16:56:06 on February 21, 2007

    Entry number 194404

    keyword=Current calibration summary (0L02 and Calorimeter runs)

    ---------------------------------------------------------
    BCM Linearity calibration  Tuesday, Feb. 20  (0800--1230)
    ---------------------------------------------------------
    The 60min BCM Linearity calibration seems to have had some problems getting
    going.  AFAICT, the good data started at 11:29; stopped at 12:03.
      Halog entry (FWIW: screen snaps are good, but render incorrectly in firefox 2.0.0)
        http://www.jlab.org/~adaq/halog/html/0702_archive/070220120532.html
      Data /usr/opdata/bcm_hallA/BcmLog_127
      CODA runs: 6631/26617
    The calibration was done from 68 uA down to 0.2 uA  (even higher starting
    point than requested).  See OPS Log entry
      http://opweb.acc.jlab.org/CSUEApps/elog02/elog_item.php?elog_id=1359349&PHPSESSID=35aa0b7ded92ad555a923283cd1d7fed
    
    ---------------------------------------------------------
    Summary of the Calorimeter runs to date.
    ---------------------------------------------------------
    Files                                  Q_cal   Q_cup
                                      uA         uCoulombs
    ---------------------------------------------------------
    --- Monday
    calorimeter.2007-02-12.08:54:00   0.6     260.9    257.5
    -? Nominal requested current:     1.0 uA
    -  BCM log file: BcmLog_110
    -  CODA runs:    6248/26223
    
    --- Tuesday
    calorimeter.2007-02-13.15:47:53   0.4     100.4     98.5
    -  BCM log file: BcmLog_111
    -  CODA runs:    
    
    --- Wednesday
    calorimeter.2007-02-14.08:23:17    --     -----     ----
    calorimeter.2007-02-14.08:36:55   0.14    -----    -----
    -  Nominal requested current:     0.2 uA
    -  BCM log file: BcmLog_114
    calorimeter.2007-02-14.08:55:07    --     -----     -----
    
    --- Thursday
    calorimeter.2007-02-15.08:25:46    --     -----     -----
    -  Nominal requested current:     0.2 uA
    -  CODA runs:    6428/26406
    -  BCM log file: BcmLog_115
    calorimeter.2007-02-15.09:02:15   0.4     141.0    141.6
    -  Nominal requested current:     0.5 uA
    -  CODA runs:    6430/26408
    -  BCM log file: BcmLog_116
    
    --- Friday
    calorimeter.2007-02-16.09:32:26   0.2      70.9     70.3
    -  Nominal requested current:     0.2 uA
    -  CODA runs:    6474/26452
    -  BCM log file: BcmLog_117
    calorimeter.2007-02-16.11:00:54   2.5     555.0    549.0
    -  Nominal requested current:     2.5 uA
    -  CODA runs:    6478/26458
    -  BCM log file: BcmLog_118
    
    -------------------------------------------------------------
    New runs.  Arne is working on the analysis
    -------------------------------------------------------------
    --- Sunday, Feb 18
    -  Nominal requested current:     0.2 uA
    -  CODA runs:    6559/26543.
    -  BCM log file: BcmLog_119
    NOTE: MCC called us halfway through the run to start the logging
          so the files are not complete.
    
    -  Nominal requested current:     1.0 uA
    -  CODA runs:    6567/26551
    -  BCM log file: BcmLog_121
    NOTE: Should be OK.  (I don't know what happened to BcmLog_120 though...)
    
    --- Monday, Feb 19
    -  Nominal requested current:     0.5 uA
    -  CODA runs:    6609/26595
    -  BCM log file: BcmLog_122
    NOTE: Beam trip for ~2 minutes during calorimeter run.
    
    -  Nominal requested current:     0.7 uA
    -  CODA runs:    6612/26598
    -  BCM log file: BcmLog_123
    NOTE: MCC reported problem with Calorimeter script.  Probably junk?
    
    --- Tuesday, Feb 20
    -  Nominal requested current:     5.0 uA
    -  CODA runs:    6626 / 26612
                     6627 / 26613  (5.0uA data at start, some 0.2 uA data at the end)
    -  BCM log file: BcmLog_124
    NOTE: Left arm CODA run stopped during the Calorimeter run due to an event
          limit.  Shift crew started a second pair after ~1min pause.  Calorimeter
          run is good.
    
    -  Nominal requested current:     0.2 uA
    -  CODA runs:    6627 / 26613
    -  BCM log file: BcmLog_124
    NOTE: A 0.2uA run was done right after the above 5.0 uA run without stopping
          CODA runs or BCM logging (hence same files).
    
    


    A copy of this log entry has been emailed to: camsonne@jlab.org